the total of three sisters ages is 39. Dina is half as old as Michelle and 3 years younger than Juliette. how old are the sisters?
Found 3 solutions by ankor@dixie-net.com, josgarithmetic, MathTherapy:
Answer by ankor@dixie-net.com(22740)

the total of three sisters ages is 39.
d + m + j = 39
Dina is half as old as Michelle
d = .5m
or multiply be 2
m = 2d
and 3 years younger than Juliette.
d = j - 3
or Add 3 to both sides
j = d + 3
how old are the sisters?
Replace m with 2d, replace j with (d+3) in the first equation
d + 2d + (d+3) = 39
4d = 39 - 3
4d = 36
d = 36/4
d = 9 yr is Dina's age
and
2(9) = 18 yrs is Michelle's age
and
9+3 = 12 yrs is Julie's age
:
:
Check: 9 + 18 + 12 = 39

Answer by josgarithmetic(39618)

Identical to question number 1173396. Simple to answer, and easily done in just one single variable.

Answer by MathTherapy(10552)


the total of three sisters ages is 39. Dina is half as old as Michelle and 3 years younger than Juliette. how old are the sisters?
Let Dina's age be D
Then Michelle's, and Juliette's are, 2D and D + 3, respectively
We then get: D + 2D + D + 3 = 39
4D = 36
Dina, or
You should now be able to find Michelle's and Juliette's ages!